Bob has just finished climbing a sheer cliff above a beach, and wants to figure out how high he climbed. All he has to use, however, is a baseball, a stopwatch, and a friend on the beach below with a long measuring tape. Bob is a pitcher and he knows that the fastest he can throw the ball is about ?0=32.9 m/s. Bob starts the stopwatch as he throws the ball (with no way to measure the ball's initial trajectory), and watches carefully. The ball rises and then falls, and after ?1=0.710 s the ball is once again level with Bob. Bob cannot see well enough to time when the ball hits the ground. Bob's friend then measures that the ball landed ?=125 m from the base of the cliff. How high up is Bob, if the ball started exactly 2 m above the edge of the cliff?
Answer #1

answer) from the figure

vo=32.9m/s

in vertical direction

vy=g*t/2=9.8*(0.71/2)=3.479m/s

from the diagram

\Theta=sin-1(3.479/32.9)=6.1o

so Vx=vcos\Theta=32.9*6.1=32.714m/s

Time to impact the ground (t1)=x/vx=125m/32.714=3.821s

so now the falling height of the ball=-vyt1+1/2gt12= -3.479m/s*3.821s+1/2*9.8*3.8212=58.2m

the height at which bob=58.2-2=56.2m

so the answer is 56.2m or 56m.